Crack Identification by Passive Electric Potential CT Method Using Piezoelectric Material

Shiro KUBO, Takahide SAKAGAMI, LI Shangqing and Daiki SHIOZAWA Department of Mechanical Engineering and Systems Graduate School of Engineering, Osaka University 2-1, Yamadaoka, Suita, Osaka 565-0871 Japan Phone: +81-6-6879-7304, Facsimile: +81-6-6879-7305 e-mail: kubo@mech.eng.osaka-u.ac.jp

The crack identification from physical measurements can be categorized as one of the domai/boundary inverse problems [1]. The present authors proposed the electric potential CT (computed tomography) method for identifying cracks in two- and threedimensional bodies [2-4]. This method can be categorized as an active method, since the electric potential distribution on the surface of cracked body is computer-processed for identifying the crack. The applicability of the method for the identification of cracks in homogeneous body and bonded dissimilar bodies was demonstrated by numerical simulations and experiments.

They also proposed a passive electric potential CT method for the crack identification using the piezoelectric material [5]. In this method the piezoelectric material is attached to a cracked body subjected to mechanical loading. Due to the direct piezoelectric effect the strain in the cracked body induces the electric potential distribution on the surface of the piezoelectric material, which can be used for the crack identification.

In the present paper the effects of crack length and crack location were investigated in the passive electric potential CT method. Finite element analyses were conducted taking account of the coupled mechano-electrical effect. Electric potential measurements were made using nanovoltmeter and surface electric potential meter for an aluminum alloy plate. The effects of the crack location and size on the electric potential observed on the surface of piezoelectric material were discussed. It was found that the cracks can be quantitatively identified using the electric potential distributions.

Keywords: Piezoelectric material, Intelligent structure, Crack, Nondestructive testing

REFERENCES

1. S. Kubo, Inverse Problems Related to the Mechanics and Fracture of Solids and Structures, JSME International Journal, Series I, 31 (1988) 157.

2. K. Ohji, S. Kubo, and T. Sakagami, (1985), Trans. Jpn. Soc. Mech. Eng., Ser.A, 51 (1985), pp.1818-1825 (in Japanese).

3. S. Kubo, T. Sakagami, and K. Ohji, Computational Mechanics ’86, (ed. by G. Yagawa, S.N. and Atluri), Springer-Verlag, Berlin, (1986), pp.V-339-344.

4. S. Kubo, T. Sakagami, and K. Ohji, Current Japanese Materials Research- Vol.8 Fracture Mechanics (ed. H. Okamura and K. Ogura), Elsevier Appl. Sci.œB!œ(BSoc. Mat. Sci., Japan, (1991), pp.235ffi0ffl254.

5. Li S.Q. S. Kubo, T. Sakagami, and Liu Z.X., Material Sci. Research Int., 6 (2000), pp.41-48.